19, deeantt equal 17.0. The sigpificance level to be saed in 946 asd the sample iue so be ukes n 73 Te table valuele) to we from the dhi-sqare ir entiom islare 40471 B. 13.181 C 39344 D. 12401 and 39 364 20. Consider the folorwing pairs of measarementa a s4 82 5 3 y13 3 7 2 9 Find the least squares A 254-994 squares estimates of β°and B.844-926 D.205-839 End of eam

please answer question 19 and 20
Answer #1

19) The table value is given by: χ2α/2,N-1 . we consider α/2 because the alternate hypothesis is that the variance is not equal to 17.0.

χ20.05/2,25-1 = χ20.025,24 = 39.364 (from the chi-square table)

Ans: C)39.364

20) The R code is given as:

x=c(8,5,4,6,2,5,3)

y=c(1,3,6,3,7,2,5)

fit=lm(y~x)

summary(fit)

The least square estimates are:

Ans: A) 8.54 and -0.994